# Weighted Birkhoff averages: Deterministic and probabilistic perspectives

### v2

@@ 1 Introduction
${\rm{WB}}_{N}\left(f\right)\left(\theta\right):=\frac{1}{{{A^{p,q}_{N}}}}\sum\limits_{n=0}^{N-1}{w_{p,q}\left({n/N}\right)f\left({{\mathscr{T}^{n}}\theta}\right)}\quad\text{or}\quad\frac{1}{N}\int_{0}^{N}{w\left({s/N}\right)f\left({{\mathscr{T}^{s}}\theta}\right)ds},$ (1.2)
@@
This weighted approach, initially introduced in [Las93a, Las93b, Las99] as a breakthrough, exhibits unexpectedly excellent convergence, known as “super convergence”, in the study of celestial mechanics as well as dynamical systems. This has been observed in practical computations [Las93a, Las93b, Las99, DSSY17, DY18, SY18, SM20, MS21, BC23, DM23, BM24, Boh24, CCGd24, MS25, SM25b, RB25, RKB25, SM25a, TL25a] and rigorously proved in theory [DSSY17, DY18, DM23, RB24, TL24a, TL24b, TL24c, TL25b] as arbitrary polynomial uniform convergence in the finite-dimensional setting. Here, we would like to mention that the original proof came from [DSSY17, DY18]These works originate from the research group of Prof. James Alan Yorke.. In particular, the uniform convergence could even be exponential [TL24a, TL24b, TL24c, TL25a, TL25b] in a universal sense, even in the infinite-dimensional setting [TL24c]. More precisely, the uniformity is for all initial values, while the universality is for almost all rotations with full Lebesgue measure or full probability measure, as well as observables that possess analyticity. We will provide later a detailed, rigorous and improved statement in Theorem 1.2. However, it is worth mentioning that under this weighted method, the $C^{\infty}_{0}([0,1])$ property of the weighting function is essential. Otherwise, for $C^{m}_{0}([0,1])$ weighting functions with finite $m$, only polynomial convergence of finite order can be achieved, as shown by the counterexamples constructed in [TL24b, TL24c, TL25b]. This means that the weighting functions initially and frequently utilized, such as $\sin^{2}(\pi x)$ [Las93a, Las93b, Las99], offer no advantages compared to (1.1) in practiceFrom the perspective of ergodic theory, which does not focus solely on the rate of numerical simulation, considering the corresponding convergence results remains independently interesting., especially since they have essentially the same computational cost as (1.1), as pointed out in [DSSY17, SM25a] among others.
However, in the classical ergodic theory, Birkhoff averages, with uniformly equal weights $1/N$, depressingly possess quite slow convergence. Given a mapping $\mathcal{T}$ on a topological space $X$ with a probability measure $\mu$ for which $\mathcal{T}$ is invariant and ergodic, the Birkhoff average of an observable $f$ defined on $X$ is defined as
${{\rm B}_{N}}\left(f\right)\left(x\right):=\frac{1}{N}\sum\limits_{n=0}^{N-1}{f\left({{\mathcal{T}^{n}}x}\right)},\;\;x\in X.$
@@
$\mathop{\lim\sup}\limits_{N\to+\infty}\epsilon_{N}^{-1}\left|{{{\rm B}_{N}}\left(f\right)\left(x\right)-\int_{X}{fd\mu}}\right|=+\infty,\;\;\text{a.e.}$
@@
This important observation is commonly referred to as “the absence of estimates for rates of convergence in the Birkhoff ergodic theorem”. Very recently, similar yet distinct statements were proved in [Ryz23, Ryz25a, Ryz25b, Ryz25c]§§§The first author expresses deep gratitude for the communications with Prof. Valery V. Ryzhikov. He pointed out that similar results can be achieved by introducing weighting to the Birkhoff averages in the general case. via elegant approaches. We would also like to mention other fundamental results on Birkhoff averages and related ergodic problems with slow convergence rates, for instance, [dR79, KP81, PUZ89, FS03, KP16, FS18, Fan19, KP19, HSY19a, HSY19b, Fan21, HLW21, Pod22, Pod24a, Pod24b, Ryz24] and the literature cited therein.
@@
$f\left(\theta\right)=g\left({\mathscr{T}\theta}\right)-g\left(\theta\right),\;\;\theta\in{\mathbb{T}^{d}}.$ (1.3)
@@
$g\left(\theta\right)=\sum\limits_{k\in{\mathbb{Z}^{d}}}{\frac{{\hat{f}\left(k\right)}}{{\exp\left({2\pi i\left\langle{k,\rho}\right\rangle}\right)-1}}\cdot\exp\left({2\pi i\left\langle{k,\theta}\right\rangle}\right)},\;\;\theta\in{\mathbb{T}^{d}}.$
With the universal Diophantine nonresonance for $\rho\in\mathbb{R}^{d}$ (see Definition 1.2), it can be proved that $g$ is well-defined on $\mathbb{T}^{d}$, whenever $f\in C^{L}(\mathbb{T}^{d})$ with $L>0$ sufficiently large. For more accurate statements, see [Kat04]. Therefore, by summing and averaging the co-homological equation (1.3), the convergence rate of ${{\rm B}_{N}}\left(f\right)\left(\theta\right)$ can be obtained explicitly as $\mathcal{O}^{\#}(N^{-1})$. On the other hand, results with slow convergence rates can also be explicitly constructed. To show that there is no analogue of the so-called Denjoy-Koksma inequality in the case of $\mathbb{T}^{2}$, counterexamples were constructed in [Yoc80, Yoc95a] using extremely Liouville rotations (see Section 1.1.1 for the definition) and obtained arbitrarily slow convergence via analytic observables. Unlike the case in (1.2), for uniform weights–that is, weights that are independent of the time scale NN–one often can only obtain a convergence rate slower than $\mathcal{O}^{\#}(N^{-1})$, even when considering the case of toral translation [FS18, Fan19].
@@
    The vast majority of existing results focus on the special case in (1.2) where $p=q=1$, and are more inclined towards qualitative exponential convergence. However, there is currently no quantitative universal exponential convergence for the infinite-dimensional case–which is more challenging than the finite-dimensional case–regardless of any values of $p$ and $q$. We will solve this important problem in Theorem 1.2.
@@
    There have been numerous studies on quasi-periodic and almost periodic cases for (1.2). However, the universal exponential convergence in the periodic case–which intuitively should be more straightforward–has yet to be established. We will supplement this result quantitatively in Theorem 1.2.
@@
Based on the aforementioned various motivations, the main contributions and organization of this paper are as follows. In Section 1.1, we review, summarize and extend the results of the weighted method in (1.2)–alternatively referred to as a weighted quasi-Monte Carlo method–from a deterministic perspective. In Section 1.1.1, we first introduce a limit-preserving theorem and its converse (Theorem 1.1), which is essentially an abstract averaging version of the Toeplitz theorem. This theorem allows us to directly employ such weighting to accelerate convergence, provided that the limit of the average exists in advance. In addition to this, we also introduce some fundamental concepts in both finite-dimensional and infinite-dimensional settings, such as the spatial structure, Diophantine nonresonance, and analyticity. Next, as one of the main results of this paper, Theorem 1.2 provides convergence rate results for (1.2) under various regularity conditions–especially analyticity–in a universal sense with full Lebesgue measure or full probability measure. These results cover quasi-periodic, almost periodic, and periodic cases. The majority of the results are presented in the form of quantitative exponential convergence and are applicable for any $p,q>0$. In particular, the results for the almost periodic and periodic cases are entirely new. These findings not only provide nearly complete theoretical guarantees for weighted fast computation in celestial mechanics and dynamical systems but also highlight the connections with ergodic theory, number theory, and harmonic analysis. As a new discovery, we introduce a novel weighting function in Theorem 1.3, which enables the weighted Birkhoff average to converge at a rate that is nearly linear-exponential. All the above results are stated in the toral translation form for brevity, as explained previously. They could also be restated in the conjugated form as those in [DSSY17, DY18, DM23, RB24], which we do not pursue here. In Section 1.1.2, we revisit at least five applications of this weighted method in physically related problems. As one of the important applications, it can compute Fourier coefficients with an exponential convergence rate. In Section 1.1.3, we rigorously derive the corresponding approximation formula in a weighted manner and, for the first time, provide a proof of the exponential convergence of the weighted average and determine the effective order, including both finite-dimensional and infinite-dimensional cases. This result, which is not a direct consequence of Theorem 1.2, significantly differs from existing results and addresses substantial challenges, thus constituting a primary contribution of this paper. Apart from the deterministic results, we also aim to explore the probabilistic perspectives in Section 1.2. Since the various weighted laws of large numbers are almost trivial at this point, we choose not to state all of them specifically. Based on historical results, Section 1.2.1 presents the weighted strong laws of large numbers (Theorems 1.6 and 1.7); Section 1.2.2 establishes the weighted central limit theorem (Theorem 1.8) and leaves a conjecture open. To clarify the new contributions of this paper and to compare them with the historical literature, we add Section 2 as a concise summary for readers. Finally, Section 3 provides the proof of the deterministic results, while Section 4 contains the proof of the probabilistic results.
@@ 1.1.1 The weighted quasi-Monte Carlo method
The weighted method presented in this paper, as illustrated in (1.2), is also a quasi-Monte Carlo technique and can be appropriately termed the (non-uniformly) weighted quasi-Monte Carlo method. Here, the term “non-uniformly” means that the weights depend on the time scale, and we would like to omit it in the following discussions for brevity. As previously mentioned, it significantly accelerates calculations involving averages (or integrals) in many dynamical systems of interest. Intuitively speaking, based on the weighted approach and the excellent compact support property of the weighting function $w_{p,q}(x)$, this weighted quasi-Monte Carlo method significantly reduces the weight of the initial and terminal data, focusing more on the intermediate data. The elaboration of the weighting procedure is shown explicitly in Figure 1. However, it is important to note that, the intermediate data is not necessarily crucial either, and the weight of a significant portion of the data can also be substantially reduced (i.e., by utilizing some new weighting functions) while still ensuring rapid convergence of arbitrary polynomial or even exponential types [TL24b]. Therefore, we can actually allow the data to be sparse.
@@
When applying the weighting function $w_{p,q}(x)$ to the original average, it is necessary to ensure that it does not change the original limit. Without this assurance, this weighted method may not be significant, particularly in terms of accelerating the convergence rate. The theorem below essentially represents an abstract averaging version of the Teoplitz theorem. Dealing with the Banach space $\mathbb{R}$ endowed with the sup-norm, it was claimed in [DSSY17] this assertion without a proof. In this paper, we provide a rigorous and concise proof for this intuitive phenomenon, as detailed in Section 3.1.
@@
$\frac{1}{N}\sum\limits_{n=1}^{N}{{a_{n}}}\to a\in\mathscr{B}\;\;\text{in the $\|\cdot\|_{\mathscr{B}}$ norm.}$ (1.4)
@@
$\frac{1}{{{A^{p,q}_{N}}}}\sum\limits_{n=1}^{N}{w_{p,q}\left({n/N}\right){a_{n}}}\to a\in\mathscr{B}\;\;\text{in the $\|\cdot\|_{\mathscr{B}}$ norm.}$ (1.5)
@@
Even if the convergence of the original average (1.4) is assumed, this weighted quasi-Monte Carlo method may not always significantly increase the convergence rate. For instance, if $a_{n}=1/n$, it is evident that (1.4) converges at a rate of $\mathcal{O}^{\#}((\log N)/N)$, while (1.5) converges at a rate of $\mathcal{O}^{\#}(1/N)$, which only eliminates the logarithmic factor. See also Section 1.1.2 for more cases.
The primary significance of Theorem 1.1 is to demonstrate that anything observed using the law of large numbers can also be observed using this weighted quasi-Monte Carlo method. Theorem 1.1 also tells us that even if the weighted average (1.5) converges, the original one (1.4) does not necessarily converge. Therefore, in numerical simulations [Las93a, Las93b, Las99, DSSY17, DY18, SY18, SM20, MS21, BC23, DM23, BM24, Boh24, CCGd24, MS25, RB25, RKB25, SM25a, SM25b, TL25a], the theoretical convergence of the average (1.4) must be established in advance (e.g., via ergodicity) before considering weighted acceleration.
We next present a theoretical result (Theorem 1.2) that demonstrates the universal arbitrary polynomial and exponential convergence for quasi-periodic, almost periodic, and periodic dynamical systems via this weighted quasi-Monte Carlo method, for all $p,q>0$. The results for quasi-periodic and almost periodic systems have been obtained under various settings in [DSSY17, Theorem 3.1], [DY18, Theorem 1.1], [DM23, Theorem 3], [TL24a, Theorem 2.1, Corollaries 2.1, 3.1 and 3.2], [TL24c, Theorem 4.4], and [TL25a, Theorems 6.1 and 6.2], mainly for p=q=1p=q=1***Here, we emphasize that the analysis for exponential convergence becomes more challenging when $p,q>0$ in general, as opposed to the simpler case where $p=q=1$.. Specifically, the result concerning the universal exponential convergence in the almost periodic case is significantly improved here, with the exponent explicitly depending on the spatial structure. The exponential convergence for the periodic case is presented here for the first time, although it can be regarded as a variation of [TL24c, Theorem 3.5] (if we ignore the quantitative aspects), and we would also like to mention the very recent work [RB24, Theorem 2.2] on arbitrary polynomial convergence for the periodic case with $C_{0}^{\infty}([0,1])$ weighting functions. For the sake of simplicity, we prefer not to present such results with extensions to $p,q>0$ in general Banach spaces here. Interested readers are referred to [TL24a, TL24b, TL24c] for further details.
@@
$\mathbb{T}_{\sigma}^{d}:=\left\{{x={{({x_{j}})}_{1\leqslant j\leqslant d}}:\;\;{x_{j}}\in\mathbb{C},\;\operatorname{Re}{x_{j}}\in\mathbb{T}^{1},\;\left|{\operatorname{Im}{x_{j}}}\right|\leqslant\sigma},\;\text{for all}\;1\leqslant j\leqslant d\right\}.$ (1.6)
@@
$f=\sum\limits_{k\in\mathbb{Z}^{d}}{\hat{f}(k)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}\quad\text{with}\quad\|f\|_{\sigma,d}:=\sum\limits_{k\in\mathbb{Z}^{d}}{\|{\hat{f}(k)}\|_{\ell^{\infty}}{\exp\left(2\pi\sigma{{\|k\|_{\ell^{1}}}}\right)}}<+\infty.$
@@
    The discrete case: for any $0\neq k\in\mathbb{Z}^{d}$ and $n\in\mathbb{Z}$, $\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma\left\|k\right\|_{{\ell^{1}}}^{-\tau}$, provided $\tau\geqslant d$;
@@
Here, $\tau$ is termed the Diophantine exponent of $\rho$. We say that $d$ in (I) and d1d-1 in (II) within Definition 1.2 are critical Diophantine exponents. This is because the set of such rotations in $\mathbb{R}^{d}$ admits zero Lebesgue measure but full Hausdorff dimension (e.g, $\sqrt{2}$ and the golden ratio $(\sqrt{5}+1)/2$ for $d=1$). They do not exist when the Diophantine exponent is smaller, as can be proved by Dirichlet’s principle. Conversely, when the Diophantine exponent is larger, the set of these rotations admits full Lebesgue measure and can thus be termed “universal” (e.g., $\pi$ for $d=1$). For further details in this aspect, see [Her79]. If a rotation is not Diophantine, we refer to it as Liouville. Further classifications exist for Liouville rotations, such as Bruno (Brjuno/Bryuno) rotations, which often essentially characterize the optimal arithmetic properties that are specific to various analytic dynamical systems in a sum/integral way [Yoc95b, Yoc02].
However, when considering the infinite-dimensional torus $\mathbb{T}^{\infty}$} [Bou05, Koz20, Koz21, MP21, BCGP24, CGP24], it becomes necessary to impose some spatial structure–which may not be unique–to prevent the Fourier series expansions from blowing up, for instance. Here, we directly employ the concepts from [Bou05], which were systematically refined in [MP21] and are primarily utilized to investigate quasi-periodic and almost periodic solutions in PDEs via KAM theory, including the nonlinear Schrödinger equation. This is because, once the corresponding toral translation conjugation theorem is established, the results of this paper can be directly applied to these PDEs.
@@
$\mathbb{Z}_{*}^{\infty}:=\left\{{k\in{\mathbb{Z}^{\infty}}:\;\;{{\left|k\right|}_{\eta}}:=\sum\limits_{j\in\mathbb{N}}{{{\left\langle j\right\rangle}^{\eta}}\left|{{k_{j}}}\right|}<+\infty,\;\;\left\langle j\right\rangle:=\max\left\{{1,\left|j\right|}\right\}}\right\}.$
@@
$\mathbb{T}_{\sigma}^{\infty}:=\left\{{x={{({x_{j}})}_{j\in\mathbb{N}}}:\;\;{x_{j}}\in\mathbb{C},\;\operatorname{Re}{x_{j}}\in\mathbb{T}^{1},\;\left|{\operatorname{Im}{x_{j}}}\right|\leqslant\sigma{{\left\langle j\right\rangle}^{\eta}}},\;\text{for all}\;j\in\mathbb{N}\right\}.$

### v3

@@ 1 Introduction
${\rm{WB}}_{N}\left(f\right)\left(\theta\right):=\frac{1}{{{A^{p,q}_{N}}}}\sum\limits_{n=0}^{N-1}{w_{p,q}\left({n/N}\right)f\left({{\mathscr{T}^{n}}\theta}\right)}\quad\text{or}\quad\frac{1}{N}\int_{0}^{N}{w_{{p,q}}\left({s/N}\right)f\left({{\mathscr{T}^{s}}\theta}\right)ds},$ (1.2)
@@
This weighted approach, initially introduced as a breakthrough, notably in [Las93a, Las93b, Las99], exhibits unexpectedly excellent convergence, known as “super convergence”, in the study of celestial mechanics as well as dynamical systems. This has been observed in practical computations [Las93a, Las93b, Las99, DSSY17, DY18, SY18, SM20, MS21, BC23, DM23, BM24, Boh24, CCGd24, BdJC25, BHS25, BBC25, GZE25, MS25, SM25b, RB25, RKB25, SM25a, TL25b, ZXH25] and rigorously proved in theory [DSSY17, DY18, DM23, RB24, TL24a, TL24b, TL25a, TL25c] as arbitrary polynomial uniform convergence in the finite-dimensional setting. Here, we would like to mention that the original proof came from [DSSY17, DY18]These works originate from the research group of Prof. James Alan Yorke.. In particular, the uniform convergence could even be exponential [TL24a, TL24b, TL25a, TL25b, TL25c] in a universal sense, even in the infinite-dimensional setting [TL24b]. More precisely, the uniformity is for all initial values, while the universality is for almost all rotations with full Lebesgue measure or full probability measure, as well as observables that possess analyticity. We will provide later a detailed, rigorous and improved statement in Theorem 1.2. However, it is worth mentioning that under this weighted method, the $C^{\infty}_{0}([0,1])$ property of the weighting function is essential. Otherwise, for $C^{m}_{0}([0,1])$ weighting functions with finite $m$, only polynomial convergence of finite order can be achieved in general, as shown by the counterexamples constructed in [TL24b, TL25a, TL25c]. This means that the weighting functions initially and frequently utilized, such as $\sin^{2}(\pi x)$ [Las93a, Las93b, Las99], offer no advantages compared to (1.1) in practiceFrom the perspective of ergodic theory, which does not focus solely on the rate of numerical simulation, considering the corresponding convergence results remains independently interesting., especially since they have essentially the same computational cost as (1.1), as pointed out in [DSSY17, SM25a] among others.
However, in the classical ergodic theory, Birkhoff averages, with uniformly equal weights $N^{-1}$, depressingly possess quite slow convergence. Given a mapping $\mathcal{T}$ on a topological space $X$ with a probability measure $\mu$ for which $\mathcal{T}$ is invariant and ergodic, the Birkhoff average of an observable $f$ defined on $X$ is defined as
${{\rm B}_{N}}\left(f\right)\left(x\right):=\frac{1}{N}\sum\limits_{n=0}^{N-1}{f\left({{\mathcal{T}^{n}}x}\right)},\quad x\in X.$
@@
$\mathop{\lim\sup}\limits_{N\to+\infty}\epsilon_{N}^{-1}\left|{{{\rm B}_{N}}\left(f\right)\left(x\right)-\int_{X}{fd\mu}}\right|=+\infty,\quad\text{a.e.}$
@@
This important observation is commonly referred to as “the absence of estimates for rates of convergence in the Birkhoff ergodic theorem”. Very recently, similar yet distinct statements were proved in [Ryz23, Ryz25a, Ryz25b, Ryz25c]§§§Z. Tong expresses deep gratitude for the communications with Prof. Valery V. Ryzhikov. He pointed out that similar results can be achieved by introducing weighting to the Birkhoff averages in the general case. via elegant approaches. We would also like to mention other fundamental results on Birkhoff averages and related ergodic problems with slow convergence rates, for instance, [dR79, KP81, PUZ89, LOT99, FS03, KT03, KP16, FS18, Fan19, KP19, HSY19a, HSY19b, Fan21, HLW21, Pod22, Pod24a, Pod24b, Ryz24] and the literature cited therein.
@@
$f\left(\theta\right)=g\left({\mathscr{T}\theta}\right)-g\left(\theta\right),\quad\theta\in{\mathbb{T}^{d}}.$ (1.3)
@@
$g\left(\theta\right)=\sum\limits_{0\neq k\in{\mathbb{Z}^{d}}}{\frac{{\widehat{f}\left(k\right)}}{{\exp\left({2\pi i\left\langle{k,\rho}\right\rangle}\right)-1}}\cdot\exp\left({2\pi i\left\langle{k,\theta}\right\rangle}\right)},\quad\theta\in{\mathbb{T}^{d}}.$
With the universal Diophantine nonresonance for $\rho\in\mathbb{R}^{d}$ (see Definition 1.2), it can be proved that $g$ is well-defined on $\mathbb{T}^{d}$, whenever $f\in C^{L}(\mathbb{T}^{d})$ with $L>0$ sufficiently large. For more accurate statements, see [Kat04]. Therefore, by summing and averaging the co-homological equation (1.3), the convergence rate of ${{\rm B}_{N}}\left(f\right)\left(\theta\right)$ can be obtained explicitly as $\mathcal{O}^{\#}(N^{-1})$. On the other hand, results with slow convergence rates can also be explicitly constructed. To show that there is no analogue of the so-called Denjoy-Koksma inequality in the case of $\mathbb{T}^{2}$, counterexamples were constructed in [Yoc80, Yoc95a] using extremely Liouville rotations (see Section 1.1.1 for the definition) and obtained arbitrarily slow convergence via analytic observables. Unlike the case in (1.2), for uniform weights—that is, weights that are independent of the time scale NN—one often can only obtain a convergence rate slower than $\mathcal{O}^{\#}(N^{-1})$, even when considering the case of toral translation [FS18, Fan19].
@@
    The vast majority of existing results focus on the special case in (1.2) where $p=q=1$, and are more inclined towards qualitative exponential convergence. However, there is currently no quantitative universal exponential convergence for the infinite-dimensional case—which is more challenging than the finite-dimensional case—regardless of any values of $p$ and $q$. We will solve this important problem in Theorem 1.2.
@@
    There have been numerous studies on quasi-periodic and almost periodic cases for (1.2). However, the universal exponential convergence in the periodic case—which intuitively should be more straightforward—has yet to be established. We will supplement this result quantitatively in Theorem 1.2.
@@
Based on the aforementioned various motivations, the main contributions and organization of this paper are as follows. In Section 1.1, we review, summarize and extend the results of the weighted method in (1.2)—alternatively referred to as a weighted quasi-Monte Carlo method—from a deterministic perspective. In Section 1.1.1, we first introduce a limit-preserving theorem and its converse (Theorem 1.1), which is essentially an abstract averaging version of the Toeplitz theorem. This theorem allows us to directly employ such weighting to accelerate convergence, provided that the limit of the average exists in advance. In addition to this, we also introduce some fundamental concepts in both finite-dimensional and infinite-dimensional settings, such as the spatial structure, Diophantine nonresonance, and analyticity. Next, as one of the main results of this paper, Theorem 1.2 provides convergence rate results for (1.2) under various regularity conditions—especially analyticity—in a universal sense with full Lebesgue measure or full probability measure. These results cover quasi-periodic, almost periodic, and periodic cases. The majority of the results are presented in the form of quantitative exponential convergence and are applicable for any $p,q>0$. In particular, the results for the almost periodic and periodic cases are entirely new. These findings not only provide nearly complete theoretical guarantees for weighted fast computation in celestial mechanics and dynamical systems but also highlight the connections with ergodic theory, number theory, and harmonic analysis. As a new discovery, we introduce a novel weighting function in Theorem 1.3, which enables the weighted Birkhoff average to converge at a rate that is nearly linear-exponential. All the above results are stated in the toral translation form for brevity, as explained previously. They could also be restated in the conjugated form as those in [DSSY17, DY18, DM23, RB24], which we do not pursue here. In Section 1.1.2, we revisit at least five applications of this weighted method in physically related problems. As one of the important applications, it can compute Fourier coefficients with an exponential convergence rate. In Section 1.1.3, we rigorously derive the corresponding approximation formula in a weighted manner and, for the first time, provide a proof of the exponential convergence of the weighted average and determine the effective order, including both finite-dimensional and infinite-dimensional cases. This result, which is not a direct consequence of Theorem 1.2, significantly differs from existing results and addresses substantial challenges, thus constituting a primary contribution of this paper. We also briefly discuss, in Section 1.1.4, which factors may affect the “practical” convergence rate, thereby providing a more dynamical-systems perspective for numerical simulations. Apart from the deterministic results, we also aim to explore the probabilistic perspectives in Section 1.2. Since the various weighted laws of large numbers are almost trivial at this point, we choose not to state all of them specifically. Based on historical results, Section 1.2.1 presents the weighted strong laws of large numbers (Theorems 1.6 and 1.7); Section 1.2.2 establishes the weighted central limit theorem (Theorem 1.8) and leaves a conjecture open. To clarify the new contributions of this paper and to compare them with the historical literature, we add Section 2 as a concise summary for readers. Finally, Section 3 provides the proof of the deterministic results, while Section 4 contains the proof of the probabilistic results.
@@ 1.1.1 The weighted quasi-Monte Carlo method
The weighted method presented in this paper, as illustrated in (1.2), is also a quasi-Monte Carlo technique and can be appropriately termed the (non-uniformly) weighted quasi-Monte Carlo method. Here, the term “non-uniformly” means that the weights depend on the time scale, and we would like to omit it in the following discussions for brevity. As previously mentioned, it significantly accelerates calculations involving averages (or integrals) in many dynamical systems of interest. Intuitively speaking, based on the weighted approach and the excellent compact support property of the weighting function $w_{p,q}(x)$, this weighted quasi-Monte Carlo method significantly reduces the weight of the initial and terminal data, focusing more on the intermediate data. The elaboration of the weighting procedure is shown explicitly in Figure 1. However, it is important to note that, the intermediate data is not necessarily crucial either, and the weight of a significant portion of the data can also be substantially reduced (i.e., by utilizing some new weighting functions) while still ensuring rapid convergence of arbitrary polynomial or even exponential types [TL25a]. Therefore, we can actually allow the data to be sparse.
@@
When applying the weighting function $w_{p,q}(x)$ to the original average, it is necessary to ensure that it does not change the original limit. Without this assurance, this weighted method may not be significant, particularly in terms of accelerating the convergence rate. The theorem below essentially represents an abstract averaging version of the Toeplitz theorem. Dealing with the Banach space $\mathbb{R}$ endowed with the sup-norm, it was claimed in [DSSY17] this assertion without a proof. In this paper, we provide a rigorous and concise proof for this intuitive phenomenon, as detailed in Section 3.1.
@@
$\frac{1}{N}\sum\limits_{n=1}^{N}{{a_{n}}}\to a\in\mathscr{B}\quad\text{in the $\|\cdot\|_{\mathscr{B}}$ norm.}$ (1.4)
@@
$\frac{1}{{{A^{p,q}_{N}}}}\sum\limits_{n=1}^{N}{w_{p,q}\left({n/N}\right){a_{n}}}\to a\in\mathscr{B}\quad\text{in the $\|\cdot\|_{\mathscr{B}}$ norm.}$ (1.5)
@@
Even if the convergence of the original average (1.4) is assumed, this weighted quasi-Monte Carlo method may not always significantly increase the convergence rate. For instance, if $a_{n}=n^{-1}$, it is evident that (1.4) converges at a rate of $\mathcal{O}^{\#}(N^{-1}\log N)$, while (1.5) converges at a rate of $\mathcal{O}^{\#}(N^{-1})$, which only eliminates the logarithmic factor. See also Section 1.1.2 for more cases.
The primary significance of Theorem 1.1 is to demonstrate that anything observed using the law of large numbers can also be observed using this weighted quasi-Monte Carlo method. Theorem 1.1 also tells us that even if the weighted average (1.5) converges, the original one (1.4) does not necessarily converge. Therefore, in numerical simulations [Las93a, Las93b, Las99, DSSY17, DY18, SY18, SM20, MS21, BC23, DM23, BM24, Boh24, CCGd24, BdJC25, BHS25, BBC25, GZE25, MS25, RB25, RKB25, SM25a, SM25b, TL25b, ZXH25], the theoretical convergence of the average (1.4) must be established in advance (e.g., via ergodicity) before considering weighted acceleration.
We next present a theoretical result (Theorem 1.2) that demonstrates the universal arbitrary polynomial and exponential convergence for quasi-periodic, almost periodic, and periodic dynamical systems via this weighted quasi-Monte Carlo method, for all $p,q>0$. The results for quasi-periodic and almost periodic systems have been obtained under various settings in [DSSY17, Theorem 3.1], [DY18, Theorem 1.1], [DM23, Theorem 3], [TL24a, Theorem 2.1, Corollaries 2.1, 3.1 and 3.2], [TL24b, Theorem 4.4], and [TL25b, Theorems 6.1 and 6.2], mainly for p=q=1p=q=1***Here, we emphasize that the analysis for exponential convergence becomes more challenging when $p,q>0$ in general, as opposed to the simpler case where $p=q=1$.. Specifically, the result concerning the universal exponential convergence in the almost periodic case is significantly improved here, with the exponent explicitly depending on the spatial structure. The exponential convergence for the periodic case is presented here for the first time, although it can be regarded as a variation of [TL24b, Theorem 3.5] (if we ignore the quantitative aspects), and we would also like to mention the very recent work [RB24, Theorem 2.2] on arbitrary polynomial convergence for the periodic case with $C_{0}^{\infty}([0,1])$ weighting functions. For the sake of simplicity, we prefer not to present such results with extensions to $p,q>0$ in general Banach spaces here. Interested readers are referred to [TL24a, TL24b, TL25a] for further details.
@@
$\mathbb{T}_{\sigma}^{d}:=\left\{{x={{({x_{j}})}_{1\leqslant j\leqslant d}}:\quad{x_{j}}\in\mathbb{C},\;\operatorname{Re}{x_{j}}\in\mathbb{T}^{1},\;\left|{\operatorname{Im}{x_{j}}}\right|\leqslant\sigma},\;\text{for all}\;1\leqslant j\leqslant d\right\}.$ (1.6)
@@
$f=\sum\limits_{k\in\mathbb{Z}^{d}}{\widehat{f}(k)\exp\left({2\pi i\left\langle{k,x}\right\rangle}\right)}\quad\text{with}\quad\|f\|_{\sigma,d}:=\sum\limits_{k\in\mathbb{Z}^{d}}{\|{\widehat{f}(k)}\|_{\ell^{\infty}}{\exp\left(2\pi\sigma{{\|k\|_{\ell^{1}}}}\right)}}<+\infty.$
@@
    The discrete case: for any $0\neq k\in\mathbb{Z}^{d}$ and any $n\in\mathbb{Z}$, $\left|{\left\langle{k,\rho}\right\rangle-n}\right|>\gamma\left\|k\right\|_{{\ell^{1}}}^{-\tau}$, provided $\tau\geqslant d$;
@@
Here, $\tau$ is termed the Diophantine exponent of $\rho$. We say that $d$ in (I) and d1d-1 in (II) within Definition 1.2 are critical Diophantine exponents. This is because the set of such rotations in $\mathbb{R}^{d}$ admits zero Lebesgue measure but full Hausdorff dimension (e.g, $\sqrt{2}$ and the golden ratio $(\sqrt{5}+1)/2$ for $d=1$). They do not exist when the Diophantine exponent is smaller, as can be proved by Dirichlet’s principle. Conversely, when the Diophantine exponent is larger, the set of these rotations admits full Lebesgue measure and can thus be termed “universal” (e.g., $\pi$ for $d=1$). For further details in this aspect, see [Her79]. If a rotation is not Diophantine, we refer to it as Liouville. Further classifications exist for Liouville rotations, such as Brjuno (Bruno/Bryuno) rotationsWe specifically consider Brjuno rotations that are strictly weaker than Diophantine in the sense of nonresonance (i.e., excluding the Diophantine case itself), thus belonging to the Liouville class., which often essentially characterize the optimal arithmetic properties that are specific to various analytic dynamical systems in a sum/integral way [Yoc95b, Yoc02].
However, when considering the infinite-dimensional torus $\mathbb{T}^{\infty}$} [Bou05, Koz20, Koz21, MP21, BCGP24, CGP24], it becomes necessary to impose some spatial structure—which may not be unique—to prevent the Fourier series expansions from blowing up, for instance. Here, we directly employ the concepts from [Bou05], which were systematically refined in [MP21] and are primarily utilized to investigate quasi-periodic and almost periodic solutions in PDEs via KAM theory, including the nonlinear Schrödinger equation. This is because, once the corresponding toral translation conjugation theorem is established, the results of this paper can be directly applied to these PDEs.
@@
$\mathbb{Z}_{*}^{\infty}:=\left\{{k\in{\mathbb{Z}^{\infty}}:\quad{{\left|k\right|}_{\eta}}:=\sum\limits_{j\in\mathbb{N}}{{{\left\langle j\right\rangle}^{\eta}}\left|{{k_{j}}}\right|}<+\infty,\;\left\langle j\right\rangle:=\max\left\{{1,\left|j\right|}\right\}}\right\}.$
@@
$\mathbb{T}_{\sigma}^{\infty}:=\left\{{x={{({x_{j}})}_{j\in\mathbb{N}}}:\quad{x_{j}}\in\mathbb{C},\;\operatorname{Re}{x_{j}}\in\mathbb{T}^{1},\;\left|{\operatorname{Im}{x_{j}}}\right|\leqslant\sigma{{\left\langle j\right\rangle}^{\eta}}},\;\text{for all}\;j\in\mathbb{N}\right\}.$